OpenStudy (anonymous):

solve this problem backwards. Gabriel went to a theme park with some money. He spent half of it on a Harry Potter wand.3/4 to buy a present for his little sister,$8.00 to buy himself a drink. And $2.00 to buy a souvenir How much did he start with?

OpenStudy (mathstudent55):

The last amount he spent was $2. Before that he spent $8. That is a total of $10 spent on the last two items. He started with x amount. He spent half of x, or 0.5x. Now he had 0.5x left. Then he spent 3/4 of what was left. That means was left with 1/4 of 0.5x. 1/4 of 0.5x = 0.25 * 0.5x = 0.125x We know from above that the amount of money he had left to spend after the first two purchases was $10. That means that 0.125x = $10 Solve for x.